# TinyGPT-V: Efficient Multimodal Large Language Model via Small Backbones

### v2

@@ Abstract
In recent years, multimodal large language models (MLLMs) such as GPT-4V have demonstrated remarkable advancements, excelling in a variety of vision-language tasks. Despite their prowess, the closed-source nature and computational demands of such models limit their accessibility and applicability. This study introduces TinyGPT-V, a novel open-source MLLM, designed for efficient training and inference across various vision-language tasks, including image captioning (IC) and visual question answering (VQA). Leveraging a compact yet powerful architecture, TinyGPT-V integrates the Phi-2 language model with pre-trained vision encoders, utilizing a unique mapping module for visual and linguistic information fusion. With a training regimen optimized for small backbones and employing a diverse dataset amalgam, TinyGPT-V requires significantly lower computational resources—24GB for training and as little as 8GB for inference—without compromising on performance. Our experiments demonstrate that TinyGPT-V, with its language model 2.8 billion parameters, achieves comparable results in VQA and image inference tasks to its larger counterparts while being uniquely suited for deployment on resource-constrained devices through innovative quantization techniques. This work not only paves the way for more accessible and efficient MLLMs but also underscores the potential of smaller, optimized models in bridging the gap between high performance and computational efficiency in real-world applications. Additionally, this paper introduces a new approach to multimodal large language models using smaller backbones. Our code and training weights are available in https://github.com/DLYuanGod/TinyGPT-V.
11footnotetext: Zhaoxu is a visiting student in the LAIR lab at Lehigh University.
@@ 1 Introduction
In recent years, the field of artificial intelligence has seen significant advancements through the development of multimodal large language models (MLLMs), such as GPT-4V, which have shown exceptional performance across a range of vision-language tasks (Yang et al., 2023). Despite GPT-4V’s impressive capabilities, its closed-source nature limits its widespread application and adaptability. In contrast, the open-source landscape for MLLMs is rapidly evolving, presenting models like LLaVA and MiniGPT-4 that excel in image captioning (IC), visual question answering (VQA) often comparable GPT-4V in these areas (Dai et al., 2023; Liu et al., 2023a; b; Zhu et al., 2023). Notably, MiniGPT-v2 (Chen et al., 2023) has demonstrated superior performance in various visual grounding and question-answering tasks. However, its training code remains proprietary, which poses challenges for community-driven advancements and adaptability.
@@
Figure 2: Comparison of TinyGPT-V with other current MLLMs models shows that TinyGPT-V achieves cost-effective, efficient, and high-performing with fewer parameters.
@@ 3.2 Training Stages
Table 1: The full list of datasets used by TinyGPT-V during training.
@@ 4.1 Training
Figure 7: Changes in loss during the training stage of TinyGPT-V.
@@ 4.2 Evaluation
Table 3: Comparison of reasoning answers from different Models. Text in red indicates incorrect suggestions. The TinyGPT-V’s answer was short and precise.

### v3

@@ Abstract
In recent years, multimodal large language models (MLLMs) such as GPT-4V have demonstrated remarkable advancements, excelling in a variety of vision-language tasks. Despite their prowess, the closed-source nature and computational demands of such models limit their accessibility and applicability. This study introduces TinyGPT-V, a novel open-source MLLM, designed for efficient training and inference across various vision-language tasks, including image captioning (IC) and visual question answering (VQA). Leveraging a compact yet powerful architecture, TinyGPT-V integrates the Phi-2 language model with pre-trained vision encoders, utilizing a unique mapping module for visual and linguistic information fusion. With a training regimen optimized for small backbones and employing a diverse dataset amalgam, TinyGPT-V requires significantly lower computational resources—24GB for training and as little as 8GB for inference—without compromising on performance. Our experiments demonstrate that TinyGPT-V, with its language model 2.8 billion parameters, achieves comparable results in VQA and image inference tasks to its larger counterparts while being uniquely suited for deployment on resource-constrained devices through innovative quantization techniques. This work not only paves the way for more accessible and efficient MLLMs but also underscores the potential of smaller, optimized models in bridging the gap between high performance and computational efficiency in real-world applications. Additionally, this paper introduces a new approach to multimodal large language models using smaller backbones. Our code and training weights are available in the supplementary material.
@@ 1 Introduction
Figure 1: Comparison of TinyGPT-V with other current MLLMs models shows TinyGPT-V achieves cost-effective, efficient, and high-performing with fewer parameters.
In recent years, the field of artificial intelligence has seen significant advancements through the development of multimodal large language models (MLLMs), such as GPT-4V, which have shown exceptional performance across a range of vision-language tasks (Yang et al., 2023). Despite GPT-4V’s impressive capabilities, its closed-source nature limits its widespread application and adaptability. In contrast, the open-source landscape for MLLMs is rapidly evolving, presenting models like LLaVA and MiniGPT-4 that excel in image captioning (IC), visual question answering (VQA) often comparable GPT-4V in these areas Dai et al. (2023); Liu et al. (2023a; b); Zhu et al. (2023). Notably, MiniGPT-v2 (Chen et al., 2023) has demonstrated superior performance in various visual grounding and question-answering tasks. However, its training code remains proprietary, which poses challenges for community-driven advancements and adaptability.
@@ 3.2 Training Stages
Table 1: The full list of datasets used by TinyGPT-MoE during training.
Figure 7: Changes in loss during the training stage of TinyGPT-V.
@@ 4.2 Evaluation
Figure 8: Comparison of reasoning answers from different Models. Text in red indicates incorrect suggestions. The TinyGPT-V’s answer was short and precise.